if 5x - x = 2x + x - 5 then x =
Answer by tutorcecilia(2152)

5x - x = 2x + x - 5 [Simplify]
4x = 3x - 5 [Set the equal equal to zero]
4x-3x = 3x-3x - 5
x = -5
.
Plug-in (x=-5) into the original equation:
5x - x = 2x + x - 5
5(-5) - (-5) = 2(-5) + (-5) - 5
-25 + 5 = -10 - 5 - 5
-20 = -20 [Checks out]